What is the density in (g/L) of Uranuim Hexafluoride (UF6)at 779mm hg and62C?

What  is the density in (g/L) of Uranuim Hexafluoride (UF6)at 779mm hg and62C?

Grade:11

1 Answers

Vikas TU
7 years ago
Learn this density formulae which is derived from the ideal gas eqn. that is
PV =nRT
and that is:
d = PM/RT
Now put the values and get the density value,
Therfore,
d = (779/760)*352/(0.0821*335) = > 12.96 g/ml.